OpenStudy (anonymous):

Find the equation of a normal line to the curve y=3√x^2-1 at the point where x=3

zepdrix (zepdrix):

\(\large\rm y=3\sqrt{x^2-1}\) Hey Jan, remember how to find slope of a `tangent line`? Finding the slope of a `normal line` is a similar process.

zepdrix (zepdrix):

We'll find the slope of the line tangent to the curve, and then the `negative reciprocal` of that value gives us the slope of our normal line.

zepdrix (zepdrix):

So umm, for tangent slope, we need a derivative. Any trouble there?

OpenStudy (anonymous):

So (x^2-1)^1/3

OpenStudy (anonymous):

Then 1/3(x^2-1)^-2/3*(2x)

zepdrix (zepdrix):

Ok great, let's clean it up a little bit.\[\large\rm y'=\frac{2x}{3(x^2-1)^{2/3}}\]Next, we want to evaluate this derivative at x=3, \(\large\rm y'(3)=?\)

OpenStudy (anonymous):

So 1

zepdrix (zepdrix):

Sorry I ran off :( The site is acting up today.. I can't stay connected for some reason. \(\large\rm y'(3)=\dfrac{2(3)}{3(3^2-1)^{2/3}}\) Hmm I think it works out to 1/2 right?